6.7 moles of carbon dioxide molecules contain how many molecules?
1
Expert's answer
2019-05-15T06:44:48-0400

One mole is the amount having the same number of particles found in 12.000 grams of 12C (which is roughly 6.02x1023). Thus, 6.7 moles of carbon dioxide contain "6.7 \cdot 6.02 \cdot 10^{23} = 4.03 \cdot 10^{24}" molecules.